darkknight:

after a vigorous soccer match,Tina and Michael decide to have a glass of their favorite refreshment. They each run in a straight line along the indicated paths at a speed of 10 ft/sec. Parametrize the motion ofTina and Michael individually. Find when and where Tina and Michael are closest to one an-other; also compute this minimum distance. Michael starts at (0,0) and goes to (200,300). Tina starts at (400,50) and goes to (-50,275). Pre-calc. problem, help!

darkknight:

I already parametrized the equations, just how do you compute the minimal distance using pre-calc. skills?

mhchen:

hmm I can't remember this, what does your parmaterized equation look like?

darkknight:

for michael x = 20t y= 30t for tina x= 400-45t y=50+22.5t

mhchen:

oh wait what was the distance formula again hmm

mhchen:

yeah square root of (x-x)^2 + (y-y)^2 something like that hmm

darkknight:

yeah, is that what I use?

mhchen:

yeah, cause we'll be trying to find when that's the smallest

darkknight:

ok, let me see if that works and i'll let you know

darkknight:

So I got 5.59016994 times the sprt of (137t^2- 1688t +5200)

mhchen:

alright cool so that looks like a quadratic I think

darkknight:

So I guess I could graph it and see what the vertex is or change the equation to vertex/standard form and find the vertex. the x-coor. should tell me when they are closest together and the y-coor. should tell me how far apart they are

darkknight:

yep, I got x= about 6.161 and y= about 3.821

mhchen:

ye

darkknight:

thanks for helping

mhchen:

np

darkknight:

What math are you in?

mhchen:

like 3 years ahead of you

darkknight:

ok then

darkknight:

hey I got another question

mhchen:

ye

darkknight:

so I have the equation ax^2+2xa^2+1=0. I should use the quadratic formula to solve for x in terms of a right. So what would my a, b, and c values be?

darkknight:

is a going to be a. b going to be 2a^2 and c going to be 1?

mhchen:

ye

darkknight:

nvr mind I got it